•Given the following set of simultaneous equations for two related goods,xandy, find the equilibrium conditions for each market if the consumers’ income isi=600. What types of goods arexandy?

•Q_dx= 22-3P_x+P_y+ 0.1i

• Q_sx=-5+15P_x

•Q_dy=32+2P_x-4P_y+0.1i

•Q_sy=-6+32P_y


1
Expert's answer
2022-03-17T14:11:29-0400

Solution


I = 600×0.1=60

Qdx=82-3px+py

Qsx=-5+15px


Qd=Qs

=18px-py=87 (equation 1)


92+2px-4py=-6+32py

=-2px+36py=98 (Equation 2)


Using equation 1 solve for py

Py=18px-87

Substitute equation 2

-2px+36(18px-87)=98

646px=3230

Px=5. Py=3


Qdx=70. Qdy=90


The two goods are substitutes

The coefficient of py is positive. Therefore if the price of good y goes up, the quantity demand for good x increases.